### v3

# On the Sylvester program and Cayley algorithm for vector partition reduction

Abstract
A vector partition problem asks for a number of nonnegative integer solutions to a system of linear Diophantine equations with integer nonnegative coefficients. J.J. Sylvester suggested to use variable elimination for the reduction of vector partition to a sum of scalar partitions. In the simplest case of two equations with positive coefficients A. Cayley performed a reduction of the corresponding double partition to a sum of scalar partitions using the algorithm subject to a set of conditions on the coefficients. We suggested a modification of the original Cayley algorithm for the cases when these conditions are not satisfied [14]. In this manuscript we generalize the modified Cayley algorithm to arbitrary number of the Diophantine equations to accomplish the Sylvester program of the vector partition reduction to a combination of scalar partitions.
Keywords: integer partition, double partition, vector partition.
2010 Mathematics Subject Classification: 11P82.
1 Introduction
The problem of integer partition into a set of integers is equivalent to counting number of nonnegative integer solutions of the Diophantine equation
$s=\sum_{i=1}^{m}x_{i}d_{i}={\bf x}\cdot{\bf d}.$ (1)
A scalar partition function $W(s,{\bf d})\equiv W(s,\{d_{1},d_{2},\ldots,d_{m}\})$ solving the above problem is a number of partitions of an integer $s$ into positive integers $\{d_{1},d_{2},\ldots,d_{m}\}$. The generating function for $W(s,{\bf d})$ has a form
$G(t,{\bf d})=\prod_{i=1}^{m}\frac{1}{1-t^{d_{i}}}=\sum_{s=0}^{\infty}W(s,{\bf d})\;t^{s}\;.$ (2)
Introducing a notation $C[t^{s}](f(t))$ for a coefficient of $t^{s}$ in the expansion of a function $f(t)$ we have
$W(s,{\bf d})=C[t^{s}]\left(\prod_{i=1}^{m}(1-t^{d_{i}})^{-1}\right).$ (3)
Sylvester proved [17] a statement about splitting of the scalar partition into periodic and non-periodic parts and showed that the partition function may be presented as a sum of ”waves”
$W(s,{\bf d})=\sum_{j=1}W_{j}(s,{\bf d})\;,$ (4)
where summation runs over all distinct factors of the elements of the generator vector ${\bf d}$. The wave $W_{j}(s,{\bf d})$ is a quasipolynomial in $s$ closely related to prime roots $\rho_{j}$ of unity being a coefficient of ${t}^{-1}$ in the series expansion in ascending powers of $t$ of a function
$F_{j}(s,t)=\sum_{\rho_{j}}\frac{\rho_{j}^{-s}e^{st}}{\prod_{k=1}^{m}\left(1-\rho_{j}^{d_{k}}e^{-d_{k}t}\right)}\;.$ (5)
The summation is made over all prime roots of unity $\rho_{j}=\exp(2\pi in/j)$ for $n$ relatively prime to $j$ (including unity) and smaller than $j$. It is possible to express any Sylvester wave $W_{j}(s,{\bf d})$ and the scalar partition $W(s,{\bf d})$ as a finite sum of the Bernoulli polynomials of higher order [11].
Consider a vector partition function $W({\bf s},{\bf D})$ counting the number of integer nonnegative solutions ${\bf x}\geq 0$ to a linear system ${\bf s}={\bf D}\cdot{\bf x}$, where ${\bf D}$ is a nonnegative integer $l\times m$ generator matrix ($l<m$). The vector partition is a natural generalization of scalar partition defined in (2-4) to the vector argument. The generating function for the vector partition reads
$G({\bf t},{\bf D})=\prod_{i=1}^{m}\frac{1}{1-{\bf t}^{{\bf c}_{i}}}=\sum_{{\bf s}}W({\bf s},{\bf D}){\bf t^{s}},\quad{\bf t^{s}}=\prod_{k=1}^{l}t_{k}^{s_{k}},\quad{\bf t}^{{\bf c}_{i}}=\prod_{k=1}^{l}t_{k}^{c_{ik}},$ (6)
where ${\bf c}_{i}=\{c_{i1},c_{i2},\ldots,c_{il}\},\ (1\leq i\leq m)$ denotes the $i$-th column of the matrix ${\bf D}=\{{\bf c}_{1},{\bf c}_{2},\ldots,{\bf c}_{m}\}$. Generalizing the coefficient notation (3) to the case of function of several variables find for $W({\bf s},{\bf D})$
$W({\bf s},{\bf D})=C[{\bf t}^{\bf s}]\left(\prod_{i=1}^{m}\frac{1}{1-{\bf t}^{{\bf c}_{i}}}\right).$ (7)
[…]
An alternative method of vector partition computation was suggested being a direct generalization of the approach developed in [11] where vector Bernoulli and Eulerian polynomials of higher order were introduced to find explicit expression for $W({\bf s},{\bf D})$ [12]. A drawback of this approach is that it does not provide any mechanism to define the chamber boundaries. As scalar partition computation requires just known functions (the Bernoulli polynomials of higher order) [11] it is promising to obtain a reduction method expressing a vector partition through a combination of scalar ones.
The first step in this direction was made in the 19th century by J.J. Sylvester [16, 17] who suggested an iterative procedure of reduction of a vector partition into a sum of scalar partitions; in [16] he gave an example of such a reduction for a double partition. Based on this approach A. Cayley developed an algorithm of double partition reduction and established a set of conditions of this method applicability [4]. Cayley showed that each column ${\bf c}_{i}$ of the two-row matrix ${\bf D}$ gives rise to a scalar partition and elements of this column must be relatively prime (we call it P-column). Another limitation of the Cayley method is that the columns ${\bf c}_{i}$ should represent noncollinear vectors. An application of the Cayley algorithm to the computation of the Gaussian polynomial coefficients is considered in [13]. A short historical excurse into the (seemingly forgotten) ideas of Sylvester and Cayley is given in Section 2 while the essence of the original Cayley algorithm is presented in Section 3.1.
The author of this manuscript considered the case when the Cayley formula for double partition fails and discussed its modification that successfully resolves the problem [14]. Specifically, the elimination of a column with the greatest common divisor (GCD) of its elements $\mbox{gcd}({\bf c}_{i})>1$ (we call it NP-column) leads not to a single scalar partition but to a weighted sum of such partitions with modified argument. This result is described in Section 3.2 and discussed in Section 3.3.
In Section 4 we extend the reduction algorithm to $l$-tuple (for arbitrary $l\geq 2$) vector partition defined by a $l\times m$ nonnegative generator matrix ${\bf D}$ made of $m$ columns ${\bf c}_{i}$. The first step of the algorithm discussed in Section 4.1 is the partial fraction expansion of the generating function (6) into $m$ terms corresponding to elimination of each of $m$ unknowns. The contribution of a column is computed in Section 4.2 and it is either a single partition for P-columns (Section 4.2.1) or a superposition of such partitions (Section 4.2.2) for NP-column. When the last element of a column is zero both the original Cayley method and its generalization cannot be applied directly so that this case requires a special treatment. In Section 4.3 we prove that the contribution of such a column to the vector partition is zero.
The case of collinear columns of the generator matrix ${\bf D}$ is addressed in Section 5 where we show that the vector partition reduces to a convolution of scalar partitions. Additional details of the iterative reduction procedure are discussed in Section 6 while the concluding remarks are presented in Section 7.
2 Sylvester-Cayley method
The problem of vector (compound) partitions has a long history that started in 1768 when Euler announced the generating function (6) for double partition ($l=2$) [7] and followed by J.J. Sylvester presenting an alternative approach in 1858. Sylvester wrote: ”Any given system of simultaneous simple equations to be solved in positive integers being proposed, the determination of the number of solutions of which they admit may in all cases be made to depend upon the like determination for one or more systems of equations of a certain fixed standard form. When a system of $r$ equations between $n$ variables of the aforesaid standard form is given, the determination of the number of solutions in positive integers of which it admits may be made to depend on the like determination for
$\frac{n(n-1)\ldots(n-r+2)}{1\cdot 2\ldots(r-1)}$
single independent equations derived from those of the given system by the ordinary process of elimination, with a slight modification; the result being obtained by taking the sum of certain numerical multiples (some positive, others negative) of the numbers corresponding to those independent determinations. This process admits of being applied in a variety of modes, the resulting sum of course remaining unaltered in value whichever mode is employed, only appearing for each such mode made up of a different set of component parts” [16].
Then he added in the footnote: ”If there be $r$ simultaneous simple equations between $n$ variables (in which the coefficients are all positive or negative integers) forming a definite system (that is, one in which no variable can become indefinitely great in the positive direction without one or more of the others becoming negative), and if the $r$ coefficients belonging to each of the same variable are exempt from a factor common to them all, and if not more than $r-1$ of the variables can be eliminated simultaneously between the $r$ equations, then the determination of the number of positive integer solutions of the given system may be made to depend on like determinations for each of $n$ derived independent systems, in each of which the number of variables and equations is one less than in the original system” [16].
In other words, Sylvester claimed that $r$-tuple vector partition for $n$ variables can be reduced to a sum of $\binom{n}{r-1}$ scalar partitions and the reduction algorithm is an iterative process based on the variable elimination. Sylvester considered a specific double partition problem as an illustration of his method and determined regions (chambers) on a plane each having a unique expression for vector partition valid in this region only. He showed that the expressions in the adjacent chambers coincide at their common boundary (see also [15]).
Two years later this approach was successfully applied by Cayley [4] to double partitions subject to some restrictions on the positive elements of the matrix ${\bf D}$ (the columns ${\bf c}_{i}$ are noncollinear, column elements should be relatively prime and for all $1\leq i\leq m$ the inequality $c_{i2}<s_{2}+2$ holds). Cayley noticed : ”The subject (as I am aware) has hardly been considered except by Professor Sylvester, and it is greatly been regretted that only an outline of his valuable researches has been published: the present paper contains the demonstration of a theorem, due to him, by which (subject to certain restrictions) the question of Double Partitions is made to depend upon the ordinary theory of Single Partitions” [4]. Cayley demonstrated that row elimination leads to a sum of $m$ scalar partition terms each corresponding to elimination of single column of the matrix ${\bf D}$.
It should be noted that both Sylvester and Cayley usually considered coefficients of the system of Diophantine equations ${\bf s}={\bf D}\cdot{\bf x}$ being positive and subject to the specific conditions, while once Cayley mentioned that the elements of the matrix ${\bf D}$ as well as $s_{i}$ “being all positive integer numbers, not excluding zero” [4] but never explicitly discussed such a case.
Both manuscripts [16] and [4] were mentioned by L.E. Dickson in the then comprehensive volume [6] on the theory of numbers devoted to Diophantine analysis, but the literature search shows that both the reduction idea of Sylvester and the result of Cayley were forgotten. There are no publications discussing Sylvester research program of vector partition reduction to scalar ones. In my opinion the ideas and results of Sylvester and Cayley deserve not to be neglected, and this manuscript is an attempt to revive them and pave way to a general algorithm for vector partition expression through scalar partitions only.
3 Cayley method for double partitions
Computation of vector partition $W({\bf s},{\bf D})$ in (7) is performed by iterative elimination of $(l-1)$ variables $t_{k},(2\leq k\leq l)$ [2]. Each elimination step includes partial fraction expansion (PFE) w.r.t. the eliminated variable with subsequent coefficient evaluation. This step is equivalent to elimination of $k$-th row of augmented $l\times(m+1)$ matrix ${\bf E}=\{{\bf s}={\bf c}_{0},{\bf c}_{1},{\bf c}_{2},\ldots,{\bf c}_{m}\}$ made of the matrix ${\bf D}$ and vector ${\bf s}$; the same time the numerical columns of ${\bf D}$ are eliminated one by one. The number of new $(l-1)\times m$ matrices is equal to $m$. This algorithm was employed in [4] for a two-row positive matrix (see also [16]). The original Cayley algorithm and its generalization are exhaustively discussed in [14] and below we present their main points.
3.1 Original Cayley algorithm
Consider the simplest vector partition case $l=2$ following algorithm described in [4]. Denote elements of matrix ${\bf E}$ columns as ${\bf s}={\bf c}_{0}=\{r,\rho\}^{T}$ and ${\bf c}_{i}=\{b_{i},\beta_{i}\}^{T},\ (1\leq i\leq m),$ where {}$\{\cdot\}^{T}$ designates transposition of a vector. Cayley specified [4] the following conditions that should be met in order to apply the algorithm. First, all fractions $b_{i}/\beta_{i}$ should be unequal, in other words, the columns ${\bf c}_{i}$ must represent noncollinear vectors. Second, the elements of each column ${\bf c}_{i}$ should be relatively prime, i.e., $\gcd(b_{i},\beta_{i})=1$. Finally, all elements of the second row should satisfy a condition: $\beta_{i}<\rho+2$.
Assuming all $b_{i},\beta_{i}>0$ perform PFE step to present $G({\bf t},{\bf D})$ as sum of $m$ fractions (${\bf t}=\{x,y\}$)
$G({\bf t},{\bf D})=\prod_{i=1}^{m}(1-x^{b_{i}}y^{\beta_{i}})^{-1}=\sum_{i=1}^{m}T_{i}(x,y),\quad T_{i}=\frac{A_{i}(x,y)}{1-x^{b_{i}}y^{\beta_{i}}},$ (8)
where the functions $A_{i}$ are rational in $x$ and rational and integral (of degree $\beta_{i}-1$) in $y$. The value of $A_{1}(x,y)$ at $y=y_{0}=x^{-b_{1}/\beta_{1}}$ reads
$A_{1}(x,y_{0})=\prod_{i\neq 1}^{m}(1-x^{b_{i}}y_{0}^{\beta_{i}})^{-1}=\prod_{i\neq 1}^{m}(1-x^{b_{i}-b_{1}\beta_{i}/\beta_{1}})^{-1}.$ (9)
Use the relation $1/(1-x)=1/(1-x^{n})\sum_{k=0}^{n-1}x^{k}$ to write $A_{1}(x,y_{0})$ as
$A_{1}(x,y_{0})=\frac{S_{1}(x,y_{0})}{\Pi_{1}(x)},\quad S_{1}(x,y_{0})=\sum_{k=0}^{\beta_{1}-1}A_{1k}(x)y_{0}^{k},\quad\Pi_{1}(x)=\prod_{i\neq 1}^{m}(1-x^{\beta_{1}b_{i}-b_{1}\beta_{i}}),$ (10)
where $A_{1k}$ are rational functions in $x$. Consider evaluation of the contribution $C_{r,\rho}^{1}$ of the term $T_{1}(x,y)$, it is equal to the coefficient
$C_{r,\rho}^{1}=C[x^{r}y^{\rho}](T_{1}(x,y))=C[x^{r}y^{\rho}]\left(\frac{A_{1}(x,y)}{1-x^{b_{1}}y^{\beta_{1}}}\right).$
Cayley shows [4] that this term evaluates to
$C_{r,\rho}^{1}$ $=$ $C[x^{r}y^{\rho}]\left(\frac{A_{1}(x,x^{-b_{1}/\beta_{1}})}{1-x^{b_{1}}y^{\beta_{1}}}\right)=C[x^{r}]\left(x^{\rho b_{1}/\beta_{1}}A_{1}(x,x^{-b_{1}/\beta_{1}})\right)$ (11)
$=$ $C[x^{r-\rho b_{1}/\beta_{1}}]\left(A_{1}(x,x^{-b_{1}/\beta_{1}})\right)=C[x^{r\beta_{1}-\rho b_{1}}]\left(A_{1}(x^{\beta_{1}},x^{-b_{1}})\right).$
Now from (9) we obtain
$A_{1}(x^{\beta_{1}},x^{-b_{1}})=\prod_{i\neq 1}^{m}(1-x^{b_{i}\beta_{1}-b_{1}\beta_{i}})^{-1},$
and arrive at
$C_{r,\rho}^{1}=C[x^{r\beta_{1}-\rho b_{1}}]\left(\prod_{i\neq 1}^{m}(1-x^{b_{i}\beta_{1}-b_{1}\beta_{i}})^{-1}\right)=C[x^{r\beta_{1}-\rho b_{1}}]\left(1/\Pi_{1}(x)\right).$ (12)
[…]
$W(\bm{s},{\bf D})=\sum_{i=1}^{m}C_{r,\rho}^{i}=\sum_{i=1}^{m}C[x^{r\beta_{i}-\rho b_{i}}]\left(\prod_{j\neq i}^{m}(1-x^{b_{j}\beta_{i}-b_{i}\beta_{j}})^{-1}\right).$
Recall the definition of the scalar partition function (3) and rewrite it as
$W(s,{\bf d})=C[t^{s}]\left(\prod_{i=1}^{m}(1-t^{d_{i}})^{-1}\right),$
to obtain an expression of the double partition as a sum of scalar partitions
$W({\bf s},{\bf D})=\sum_{i=1}^{m}W^{2}_{i}({\bf s})=\sum_{i=1}^{m}W(L_{i}({\bf s}),{\bf d}_{i}),\quad L_{i}({\bf s})=r\beta_{i}-b_{i}\rho,\quad d_{ij}=b_{j}\beta_{i}-b_{i}\beta_{j},\ j\neq i.$ (13)
This compact expression is the main result of the original Cayley algorithm presented in [4]. Introducing $2\times 2$ matrices made of the columns of augmented matrix ${\bf E}$
${\bf D}_{i0}=\{{\bf c}_{0}={\bf s},{\bf c}_{i}\},\quad{\bf D}_{ij}=\{{\bf c}_{j},{\bf c}_{i}\},\quad j\neq i,$ (14)
we observe that $L_{i}({\bf s})$ and elements $d_{ij}$ of ${\bf d}_{i}$ are given by determinants of ${\bf D}_{i0}$ and ${\bf D}_{ij}$, respectively. Noncollinearity of the columns ${\bf c}_{i}$ implies that all elements in the generator sets ${\bf d}_{i}$ are nonzero, but some of them might be negative (say, $d_{ij_{k}}<0$ for $1\leq k\leq K_{i}$). Noting that $(1-t^{-a})^{-1}=-t^{a}(1-t^{a})^{-1}$ we find an equivalent scalar partition with positive generators only
$W(L_{i}({\bf s}),{\bf d}_{i})=(-1)^{K}W(L_{i}({\bf s})+\sum_{k=1}^{K_{i}}d_{ij_{k}},|{\bf d}_{i}|),\quad|{\bf d}_{i}|=\{|d_{ij}|\}.$ (15)
The partition $W(L_{i}({\bf s})+\sum_{k=1}^{K_{i}}d_{ij_{k}},|{\bf d}_{i}|)$ in the r.h.s. of (15) corresponds to the Diophantine equation
$\sum_{j\neq i}^{m}x_{j}|d_{ij}|=L^{\prime}_{i}({\bf s})=s_{1}\beta_{i}-s_{2}b_{i}+\sum_{k=1}^{K_{i}}d_{ij_{k}},$
where $K_{i}$ integers $d_{ij_{k}}<0$ are negative and $x_{j}\geq 0$ are nonnegative integers. When $L^{\prime}_{i}({\bf s})$ is negative this equation has no solutions and we can rewrite (15) as
$W^{2}_{i}({\bf s})=W(L_{i}({\bf s}),{\bf d}_{i})=A_{i}W(L^{\prime}_{i}({\bf s}),|{\bf d}_{i}|),\quad A_{i}=(-1)^{K_{i}}H(L^{\prime}_{i}({\bf s})),$ (16)
where $H(x)$ is the Heaviside step function defined as
$H(x)=\left\{\begin{array}[]{cc}0,&x<0,\\ 1,&x\geq 0,\end{array}\right.$
and this factor determines the chamber to which the scalar partition $W^{2}_{i}({\bf s})$ contributes.
The solution (13) is not unique as we can apply the Cayley method to the matrix ${\bf E}$ with the reverse order of the rows. It corresponds to the elimination of the first row of ${\bf E}$ and gives
$W({\bf s},{\bf D})=\sum_{i=1}^{m}W(\tilde{L}_{i}({\bf s}),\tilde{\bf d}_{i}),\quad\tilde{L}_{i}({\bf s})=-L_{i}({\bf s}),\quad\tilde{\bf d}=-{\bf d}.$ (17)
It should be noted that $\tilde{L}^{\prime}_{i}({\bf s})\neq-L^{\prime}_{i}({\bf s})$ and the individual contributions $W(L_{i}({\bf s}),{\bf d}_{i})$ and $W(\tilde{L}_{i}({\bf s}),\tilde{\bf d}_{i})$ differ. This means that the order of the rows of ${\bf E}$ does not influence the value of the vector partition $W({\bf s},{\bf D})$ but it does affect individual summands $W^{2}_{i}({\bf s})$ in (16).
It is worth to repeat after Dickson [6] that the Cayley algorithm corresponds to the elimination of each variable $x_{i}$ from the system ${\bf s}={\bf D}\cdot{\bf x}$ producing $m$ equations $L_{i}={\bf d}_{i}\cdot\bm{x}^{\prime}_{i}$ where $\bm{x}^{\prime}_{i}=\{x_{1},x_{2},\ldots,x_{i-1},x_{i+1},,\ldots,x_{m}\}$ is obtained by dropping $x_{i}$ from ${\bf x}$.
3.2 Cayley method generalization
When the Cayley method fails we suggested an alternative approach [14] based on explicit computation of $A_{1}(x,y)$. Turning to $S_{1}(x,y_{0})$ in (10) write it as
$S_{1}(x,y_{0})=\prod_{i\neq 1}^{m}\frac{(1-x^{\beta_{1}b_{i}}y_{0}^{\beta_{1}\beta_{i}})}{(1-x^{b_{i}}y_{0}^{\beta_{i}})}=\prod_{i\neq 1}^{m}\left(\sum_{k_{i}=0}^{\beta_{1}-1}x^{k_{i}b_{i}}y_{0}^{k_{i}\beta_{i}}\right).$ (18)
Introduce $(m-1)$-dimensional vectors
$\bm{b}^{\prime}_{1}=\{b_{2},b_{3},\ldots,b_{m}\},\ \bm{\beta}^{\prime}_{1}=\{\beta_{2},\beta_{3},\ldots,\beta_{m}\},\ \bm{K}^{\prime}_{p}=\{k_{2},k_{3},\ldots,k_{m}\},$ (19)
with $10\leq k_{i}\leq\beta_{1}-$ 1. Expanding the r.h.s. of (18) we obtain
$S_{1}(x,y_{0})=\sum_{p=1}^{n_{1}}x^{\bm{K}^{\prime}_{p}\cdot\bm{b}^{\prime}_{1}}y_{0}^{\bm{K}^{\prime}_{p}\cdot\bm{\beta}^{\prime}_{1}},\quad n_{1}=\beta_{1}^{m-1}.$ (20)
For each $\bm{K}^{\prime}_{p}$ we have
$x^{\bm{K}^{\prime}_{p}\cdot\bm{b}^{\prime}_{1}}y_{0}^{\bm{K}^{\prime}_{p}\cdot\bm{\beta}^{\prime}_{1}}=x^{\bm{K}^{\prime}_{p}\cdot\bm{b}^{\prime}_{1}-b_{1}(\bm{K}^{\prime}_{p}\cdot\bm{\beta}^{\prime}_{1})/\beta_{1}}=x^{\bm{K}^{\prime}_{p}\cdot\bm{b}^{\prime}_{1}-b_{1}\lfloor(\bm{K}^{\prime}_{p}\cdot\bm{\beta}^{\prime}_{1})/\beta_{1}\rfloor}y_{0}^{(\bm{K}^{\prime}_{p}\cdot\bm{\beta}^{\prime}_{1})\bmod{\beta_{1}}}\;,$
where $\lfloor a\rfloor$ denotes the greatest integer less than or equal to real number $a$. We have for $A_{1}(x,y)$
$A_{1}(x,y)=\frac{1}{\Pi_{1}(x)}\sum_{k=0}^{\beta_{1}-1}A_{1,j_{y}}(x)y^{j_{y}}\;,\quad j_{y}=(\bm{K}^{\prime}_{p}\cdot\bm{\beta}^{\prime}_{1})\bmod{\beta_{1}},$ (21)
with
$A_{1,j_{y}}(x)=\sum_{j_{x}=N_{1}^{-}}^{N_{1}^{+}}a_{1,j_{x},j_{y}}x^{j_{x}},\quad N_{1}^{-}=\min(\bm{K}^{\prime}_{p}\cdot\bm{b}^{\prime}_{1}-b_{1}\lfloor(\bm{K}^{\prime}_{p}\cdot\bm{\beta}^{\prime}_{1})/\beta_{1}\rfloor),\quad N_{1}^{+}=(\beta_{1}-1)|\bm{b}^{\prime}_{1}|,$ (22)
where we introduce $L_{1}$-norm of $m$-component vector $|\bm{a}|=\sum_{i=1}^{m}a_{i}$. The integer coefficients $a_{1,j_{x},j_{y}}$ are computed from the relation
$\sum_{j_{y}=0}^{\beta_{1}-1}\sum_{j_{x}=N_{1}^{-}}^{N_{1}^{+}}a_{1,j_{x},j_{y}}x^{j_{x}}y^{j_{y}}=\sum_{p=1}^{n_{1}}x^{\bm{K}^{\prime}_{p}\cdot\bm{b}^{\prime}_{1}-b_{1}\lfloor(\bm{K}^{\prime}_{p}\cdot\bm{\beta}^{\prime}_{1})/\beta_{1}\rfloor}y^{(\bm{K}^{\prime}_{p}\cdot\bm{\beta}^{\prime}_{1})\bmod{\beta_{1}}},\quad n_{1}=\beta_{1}^{m-1}.$ (23)
Using $A_{1,j_{y}}(x)$ in expression of $T_{1}$ and evaluating $C_{r,\rho}^{1}$ we obtain
$C^{1}_{r,\rho}=\bar{W}^{2}_{1}({\bf s})=\sum_{j_{x}=N_{1}^{-}}^{N_{1}^{+}}a_{1,j_{x},j_{y}}W(r-j_{x}-(\rho-j_{y})b_{1}/\beta_{1},{\bf d}_{1}),$ (24)
$j_{y}=\rho\bmod{\beta_{1}},\quad N_{1}^{+}=(\beta_{1}-1)|\bm{b}^{\prime}_{1}|,\quad d_{1j}=b_{j}\beta_{1}-b_{1}\beta_{j},\ j\neq 1.$
It is shown in [14] that the term $\bar{W}^{2}_{i}$ can be written as a weighted sum of $W_{i}^{2}$ with shifted argument
$\bar{W}_{i}^{2}({\bf s})=\sum_{{\bf j}={\bf N}_{i}^{-}}^{{\bf N}_{i}^{+}}a_{i,{\bf j}}\;\delta_{j_{y},\rho\bmod{\beta_{i}}}W_{i}^{2}(({\bf s}-{\bf j})/\beta_{i}),\quad{\bf j}=\{j_{x},j_{y}\}^{T},$ (25)
${\bf N}_{i}^{-}=\{\min(\bm{K}^{\prime}_{p}\cdot\bm{b}^{\prime}_{i}-b_{i}\lfloor(\bm{K}^{\prime}_{p}\cdot\bm{\beta}^{\prime}_{i})/\beta_{i}\rfloor),0\}^{T},\quad{\bf N}_{i}^{+}=(\beta_{i}-1)\{|\bm{b}^{\prime}_{i}|,1\}^{T},$
where the vectors $\bm{b}^{\prime}_{i},\bm{\beta}^{\prime}_{i}$ are given by
$\bm{b}^{\prime}_{i}=\{b_{1},b_{2},\ldots,b_{i-1},b_{i+1},\ldots,b_{m}\},\quad\bm{\beta}^{\prime}_{i}=\{\beta_{1},\beta_{2},\ldots,\beta_{i-1},\beta_{i+1},\ldots,\beta_{m}\}.$
Using (16) for each summand in (25) with ${\bf j}=\{j_{1},j_{2}\}^{T}$ we can rewrite this expression as
$\bar{W}_{i}^{2}({\bf s})=\sum_{{\bf j}={\bf N}_{i}^{-}}^{{\bf N}_{i}^{+}}A_{i,{\bf j}}W_{i}^{2}(({\bf s}-{\bf j})/\beta_{i}),\ \ A_{i,{\bf j}}=(-1)^{K_{i}}a_{i,{\bf j}}H(L^{\prime}_{i}(({\bf s}-{\bf j})/\beta_{i}))\delta_{j_{2},s_{2}\bmod{\beta_{i}}}.$ (26)
3.3 Cayley method for P- and NP-columns
The original Cayley reduction method presented in Section 3.1 applied to double partition with the matrix ${\bf D}$ having $m$ columns generates a sum (13) of $m$ scalar partitions. This approach fails when at least one of the columns, say the first column ${\bf c}_{1}$, has GCD of its elements larger than unity $\gcd({\bf c}_{1})=\gcd(b_{1},\beta_{1})=g_{1}>1,$ and we write $b_{1}=g_{1}b^{\star},\ \beta_{1}=g_{1}\beta^{\star}$. This case should be treated as discussed above in Section 3.2 leading to $C_{r,\rho}^{1}$ in (24).
Consider (24) and note that when $\gcd(b_{1},\beta_{1})=g_{1}=1$ the symbolic argument in scalar partitions in the summand is an integer only if the condition $j_{y}=\rho\bmod{\beta_{1}}$ holds. This means that when ${\bf c}_{1}$ is a P-column the condition $j_{y}=\rho\bmod{\beta_{1}}$ might be dropped and all terms in the sum in (24) contribute to $C_{r,\rho}^{1}$ that has an alternative compact form used in (13).
In case of NP-column ${\bf c}_{1}$ with $g_{1}>1$ we have $(\rho-j_{y})b_{1}/\beta_{1}=(\rho-j_{y})b^{\star}/\beta^{\star},$ and for given value of $\rho$ there exist exactly $g_{1}$ values of $j_{y}$ for which the scalar partition symbolic argument $r-j_{x}-(\rho-j_{y})b_{1}/\beta_{1}$ has integer value. The condition $j_{y}=\rho\bmod{\beta_{1}}$ selects only one of $g_{1}$ values of $j_{y}$ that should be used for the estimate of the contribution to the vector partition. Thus not all (but each $g_{1}$-th only) summands with integer scalar partition symbolic argument contribute and this fact prevents the compactification of the superposition (24) into a single term. An example of NP-column reduction in double partition is discussed in Appendix A.
Note that for $\beta_{1}=0$ the expression for $S_{1}(x,y_{0})$ in (18) vanishes implying vanishing contribution $C^{1}_{r,\rho}=0$. We also observe that the result (15) with $i=1$ and $\beta_{1}=0$ formally produces all $d_{1j}<0$ and $L_{1}({\bf s})=-s_{2}b_{1}$ leading to negative $L^{\prime}_{1}({\bf s})=-s_{2}b_{1}+\sum_{j=2}^{m}d_{1j}$ for any positive $s_{2}$. This means that the scalar partition $W(L_{1}({\bf s}),{\bf d}_{1})=W(L^{\prime}_{1}({\bf s}),|{\bf d}_{1}|)$ equals zero. Thus the rule of thumb for double partitions with nonnegative generators $b_{i},\beta_{i}$ and $r,\rho\geq 0$ is – the contribution of the column with $\beta_{i}=0$ is zero.
The same time we have to recall that in the course of the derivation of (15) we explicitly assumed that $\beta_{1}\neq 0$. Nevertheless it is possible to show that in case $\beta_{1}=0$ the contribution $C^{1}_{r,\rho}$ indeed vanishes. It appears to be a particular case of the general result valid for a $l$-tuple partition ($l\geq 2$) with the zero last element $c_{1l}=0$ in the first column ${\bf c}_{1}$ of the $l\times m$ matrix ${\bf D}$. The reader is referred to Section 4.3 where the details of the derivation are presented.
4 $l$-tuple partition reduction
Consider a partition $W({\bf s},{\bf D})$ for a non-negative $l\times m\ (l<m)$ matrix ${\bf D}=\{{\bf c}_{1},{\bf c}_{2},\ldots,{\bf c}_{m}\}$ made of $m$ columns ${\bf c}_{i}=\{c_{i1},c_{i2},\ldots,c_{il}\}^{T}$ in assumption of column noncollinearity. We would also use the equivalent notation $W({\bf E})\equiv W({\bf s},{\bf D})$ for the partition with corresponding augmented matrix ${\bf E}=\{{\bf c}_{0}={\bf s},{\bf c}_{1},{\bf c}_{2},\ldots,{\bf c}_{m}\}$ obtained by adding the symbolic argument column ${\bf s}$ to matrix ${\bf D}$.
4.1 Partial fraction expansion
Perform a PFE step (elimination of the last $l$-th row of ${\bf D}$) in assumption that all elements of this row are positive to present $G({\bf t},{\bf D})$ with ${\bf t}=\{t_{1},t_{2},\ldots,t_{l}\}$ as sum of $m$ fractions using additional shortcut notations
$\hat{\bf t}=\{t_{1},\ldots,t_{l-1}\},\quad\hat{\bf c}_{i}=\{c_{i1},\ldots,c_{i,l-1}\},\quad\hat{\bf s}=\{s_{1},\ldots,s_{l-1}\},\quad\hat{\bf j}=\{j_{1},\ldots,j_{l-1}\}.$
We have
$G({\bf t},{\bf D})=\prod_{i=1}^{m}(1-\hat{\bf t}^{\hat{\bf c}_{i}}t_{l}^{c_{il}})^{-1}=\sum_{i=1}^{m}T_{i}(\hat{\bf t},t_{l}),\quad T_{i}=\frac{A_{i}(\hat{\bf t},t_{l})}{1-\hat{\bf t}^{\hat{\bf c}_{i}}t_{l}^{c_{il}}},$ (27)
[…]
$A_{1}(\hat{\bf t},t_{l})=\prod_{i\neq 1}^{m}(1-\hat{\bf t}^{\hat{\bf c}_{i}}t_{l}^{c_{il}})^{-1},$ (28)
that is
$A_{1}(\hat{\bf t},\hat{\bf t}^{-\hat{\bf c}_{1}/c_{1l}})=\prod_{i\neq 1}^{m}(1-\hat{\bf t}^{\hat{\bf c}_{i}-c_{il}\hat{\bf c}_{1}/c_{1l}})^{-1}.$ (29)
Following the approach of Section 3.1 write (29) as
$A_{1}(\hat{\bf t},\hat{\bf t}^{-\hat{\bf c}_{1}/c_{1l}})=S_{1}(\hat{\bf t},t_{l0})/\Pi_{1}(\hat{\bf t}),\quad\Pi_{1}(\hat{\bf t})=\prod_{i\neq 1}^{m}(1-\hat{\bf t}^{c_{1l}\hat{\bf c}_{i}}t_{l0}^{c_{1l}c_{il}})=\prod_{i\neq 1}^{m}(1-\hat{\bf t}^{c_{1l}\hat{\bf c}_{i}-c_{il}\hat{\bf c}_{1}}).$ (30)
The numerator reads
$S_{1}(\hat{\bf t},t_{l0})=\prod_{i\neq 1}^{m}\left(\sum_{k_{i}=0}^{c_{1l}-1}\hat{\bf t}^{k_{i}\hat{\bf c}_{i}}t_{l0}^{k_{i}c_{il}}\right)=\sum_{j_{l}=0}^{c_{1l}-1}A_{1,j_{l}}(\hat{\bf t})t_{l0}^{j_{l}},$ (31)
[…]
$S_{1}(\hat{\bf t},t_{l0})=\sum_{p=1}^{n_{1}}t_{l0}^{\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1l}}\prod_{k=1}^{l-1}t_{k}^{\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1k}},\quad n_{1}=c_{1l}^{m-1},\quad t_{l0}=\hat{\bf t}^{-\hat{\bf c}_{1}/c_{1l}}.$ (32)
For each $\bm{K}^{\prime}_{p}$ we have
$t_{l0}^{\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1l}}=\prod_{k=1}^{l-1}t_{k}^{-c_{1k}(\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1k})/c_{1l}}$
and obtain
$t_{l0}^{\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1l}}\prod_{k=1}^{l-1}t_{k}^{\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1k}}=\prod_{k=1}^{l-1}t_{k}^{\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1k}-c_{1k}(\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1l})/c_{1l}}=\prod_{k=1}^{l-1}t_{k}^{\nu_{k}},$
where the rational exponent $\nu_{k}$ of $t_{k}$ is split into an integer $j_{k}$ and a fractional part. The fractional part of $\nu_{k}$ gives an exponent $j_{l}$ of $t_{l0}$, while $t_{k}^{j_{k}}$ contributes to $A_{1,j_{l}}(\hat{\bf t})$ in (31) as a factor in $\hat{\bf t}^{\hat{\bf j}}$. We observe that
$j_{l}=(\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1l})\bmod{c_{1l}},$ (33)
and obtain
[…]
This leads to
$t_{l0}^{\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1l}}\prod_{k=1}^{l-1}t_{k}^{\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1k}}=t_{l0}^{j_{l}}\hat{\bf t}^{\hat{\bf j}}=t_{l0}^{(\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1l})\bmod{c_{1l}}}\prod_{k=1}^{l-1}t_{k}^{j_{k}},$ (35)
with
$j_{k}=\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1k}-c_{1k}\lfloor(\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1l})/c_{1l}\rfloor=\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1k}-c_{1k}t,\quad 1\leq k\leq l-1.$ (36)
We arrive at the relation
$\sum_{j_{l}=0}^{c_{1l}-1}A_{1,j_{l}}(\hat{\bf t})t_{l0}^{j_{l}}=\sum_{p=1}^{n_{1}}t_{l0}^{(\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1l})\bmod{c_{1l}}}\prod_{k=1}^{l-1}t_{k}^{\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1k}-c_{1k}\lfloor(\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1l})/c_{1l}\rfloor},$ (37)
that allows to write for $A_{1,j_{l}}(\hat{\bf t})$
$A_{1,j_{l}}(\hat{\bf t})=\sum_{\hat{\bf j}=\hat{\bf N}_{1}^{-}}^{\hat{\bf N}_{1}^{+}}a_{1,\hat{\bf j},j_{l}}\hat{\bf t}^{\hat{\bf j}},\quad\hat{\bf N}_{1}^{\pm}=\{N_{11}^{\pm},N_{21}^{\pm},\ldots,N_{l-1,1}^{\pm}\}^{T},$ (38)
with
$N_{k1}^{-}=\min(\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1k}-c_{1k}\lfloor(\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1l})/c_{1l}\rfloor),\quad N_{k1}^{+}=(c_{1l}-1)|\bm{r}^{\prime}_{k1}|,\quad 1\leq k\leq l-1.$
The expansion coefficients $a_{1,\hat{\bf j},j_{l}}$ in (38) are computed directly from (37). The relation (27) leads to
$T_{1}(\hat{\bf t},t_{l})=\frac{A_{1}(\hat{\bf t},t_{l})}{1-\hat{\bf t}^{\hat{\bf c}_{i}}t_{l}^{c_{il}}}=\frac{1}{\Pi_{1}(\hat{\bf t})(1-\hat{\bf t}^{\hat{\bf c}_{1}}t_{l}^{c_{1l}})}\sum_{j_{l}=0}^{c_{1l}-1}A_{1,j_{l}}(\hat{\bf t})t_{l}^{j_{l}}\;.$ (39)
Substitute (38) in (39) to arrive at
$T_{1}({\bf t})=\frac{1}{\Pi_{1}(\hat{\bf t})(1-{\bf t}^{{\bf c}_{1}})}\sum_{{\bf j}={\bf N}_{1}^{-}}^{{\bf N}_{1}^{+}}a_{1,{\bf j}}{\bf t}^{\bf j},$ (40)
where ${\bf N}_{1}^{\pm}$ are obtained by adding $N_{l1}^{\pm}$ to $\hat{\bf N}_{1}^{\pm}$ with $N_{l1}^{-}=0,$ and $N_{l1}^{+}=(c_{1l}-1)|\bm{r}^{\prime}_{l1}|$.
4.2 Contribution evaluation
Use the definition (7) to compute the contribution $C_{\bf s}^{1}=C[{\bf t}^{\bf s}](T_{1}({\bf t}))$ of the term $T_{1}$ defined in (39)
$C_{\bf s}^{1}=C[\hat{\bf t}^{\hat{\bf s}}]\left(\Pi_{1}^{-1}(\hat{\bf t})C[t_{l}^{s_{l}}]\left(\frac{1}{(1-\hat{\bf t}^{\hat{\bf c}_{1}}t_{l}^{c_{1l}})}\sum_{j_{l}=0}^{c_{1l}-1}A_{1,j_{l}}(\hat{\bf t})t_{l}^{j_{l}}\right)\right).$ (41)
First consider the inner term
$C_{l}^{1}=C[t_{l}^{s_{l}}]\left(\frac{\sum_{j_{l}=0}^{c_{1l}-1}A_{1,j_{l}}(\hat{\bf t})t_{l}^{j_{l}}}{(1-\hat{\bf t}^{\hat{\bf c}_{1}}t_{l}^{c_{1l}})}\right)=\sum_{p=0}^{\infty}\sum_{j_{l}=0}^{c_{1l}-1}A_{1,j_{l}}(\hat{\bf t})\hat{\bf t}^{p\hat{\bf c}_{1}}C[t_{l}^{s_{l}}]\left(t_{l}^{pc_{1l}+j_{l}}\right),$ (42)
and observe that $C[t_{l}^{s_{l}}]\left(t_{l}^{pc_{1l}+j_{l}}\right)=1$ only for $pc_{1l}+j_{l}=s_{l}$ otherwise it vanishes. Find $p=(s_{l}-j_{l})/c_{1l}$ and use (31) to obtain
$C_{l}^{1}=\sum_{j_{l}=0}^{c_{1l}-1}A_{1,j_{l}}(\hat{\bf t})\hat{\bf t}^{(s_{l}-j_{l})\hat{\bf c}_{1}/c_{1l}}=\hat{\bf t}^{s_{l}\hat{\bf c}_{1}/c_{1l}}\sum_{j_{l}=0}^{c_{1l}-1}A_{1,j_{l}}(\hat{\bf t})t_{l0}^{j_{l}}=\hat{\bf t}^{s_{l}\hat{\bf c}_{1}/c_{1l}}S_{1}(\hat{\bf t},t_{l0}).$ (43)
4.2.1 Compact form
Substitute (43) in (41) to find
$C_{\bf s}^{1}=C[\hat{\bf t}^{\hat{\bf s}}]\left(\Pi_{1}^{-1}(\hat{\bf t})\hat{\bf t}^{s_{l}\hat{\bf c}_{1}/c_{1l}}S_{1}(\hat{\bf t},t_{l0})\right)=C[\hat{\bf t}^{\hat{\bf s}-s_{l}\hat{\bf c}_{1}/c_{1l}}]\left(\Pi_{1}^{-1}(\hat{\bf t})S_{1}(\hat{\bf t},t_{l0})\right)=C[\hat{\bf t}^{\hat{\bf s}-s_{l}\hat{\bf c}_{1}/c_{1l}}]\left(A_{1}(\hat{\bf t},t_{l0})\right).$
Recall that $t_{l0}=\hat{\bf t}^{-\hat{\bf c}_{1}/c_{1l}}$ and using (29,30) find
$C_{\bf s}^{1}=C[\hat{\bf t}^{\hat{\bf s}-s_{l}\hat{\bf c}_{1}/c_{1l}}]\left(A_{1}(\hat{\bf t},\hat{\bf t}^{-\hat{\bf c}_{1}/c_{1l}})\right)=C[\hat{\bf t}^{c_{1l}\hat{\bf s}-s_{l}\hat{\bf c}_{1}}]\left(A_{1}(\hat{\bf t}^{c_{1l}},\hat{\bf t}^{-\hat{\bf c}_{1}})\right)=C[\hat{\bf t}^{c_{1l}\hat{\bf s}-s_{l}\hat{\bf c}_{1}}]\left(1/\Pi_{1}(\hat{\bf t})\right),$ (44)
Employing expression (7) for $l$-tuple partition $W({\bf s},{\bf D})$ we observe
[…]
We observe that the expressions for ${\bf s}^{1}$ and ${\bf c}^{1}_{j}$ follow the same pattern and are just $2\times 2$ determinants similar to those introduced in (14). The determinants are computed using the columns ${\bf c}^{1}$ and corresponding columns of the augmented matrix ${\bf E}_{1}$ introduced in the beginning of this Section.
It should be reminded here that the final goal of the Sylvester program is the vector partition reduction to a set of scalar partitions and thus the elimination of a single (last) row of $l\times m$ matrix is just the first step of the procedure. We discuss the computational aspects of the Sylvester program further in Section 6 but now we have to underline that the reduction procedure described above is valid only when the matrix under reduction has its last row void of nonpositive elements. It follows from (45) that some columns ${\bf c}^{1}_{j}$ might have negative elements $c^{1}_{jl}<0$ in the last position.
A simple procedure similar to one leading to (15) allows to resolve this issue. Namely, apply the relation
$\left(1-{\bf t}^{{\bf c}}\right)^{-1}=-{\bf t}^{-{\bf c}}\left(1-{\bf t}^{-{\bf c}}\right)^{-1},$ (46)
to all $K_{1}$ columns ${\bf c}^{1}_{j}$ with $c^{1}_{jl}<0$ to update the zeroth column ${\bf s}^{1}$ and the partition itself:
$\tilde{\bf c}^{1}_{j}=-{\bf c}^{1}_{j},\quad\tilde{\bf s}^{1}={\bf s}^{1}+\sum_{j=1}^{K_{1}}{\bf c}^{1}_{j},\quad W(\{{\bf s}^{1},{\bf c}^{1}_{1},\ldots,{\bf c}^{1}_{m}\})=(-1)^{K_{1}}H(\tilde{s}^{1}_{l-1})W(\{\tilde{\bf s}^{1},\tilde{\bf c}^{1}_{1},\ldots,\tilde{\bf c}^{1}_{m}\}),$ (47)
where $\tilde{\bf c}^{1}_{i}$ for the columns not affected by the procedure is equal to its original value $\tilde{\bf c}^{1}_{i}={\bf c}^{1}_{i}$. The factor $H(\tilde{s}^{1}_{l-1})$ reflects the fact that the $(l-1)$-th element of the column $\tilde{\bf s}^{1}$ should be nonnegative $\tilde{s}^{1}_{l-1}\geq 0$.
Note that (43) does not work when GCD of the column ${\bf c}_{1}$ elements is larger than unity: $\gcd({\bf c}_{1})\equiv\gcd(c_{11},c_{12},\ldots,c_{1l})=g_{1}>1$. Then we have for such NP-column ${\bf c}_{1}=g_{1}{\bf c}^{\star}$ leading to $\hat{\bf c}_{1}/c_{1l}=\hat{\bf c}^{\star}/c_{l}^{\star}$ and $t_{l0}^{j_{l}}=\hat{\bf t}^{-j_{l}\hat{\bf c}_{1}/c_{1l}}=\hat{\bf t}^{-j_{l}\hat{\bf c}^{\star}/c_{l}^{\star}}$ enters the sum in (43) where $11\leq j_{l}\leq c_{1l}-$ 1, but $c_{1l}-1>c_{l}^{\star}$. The summation does not evaluate into $S_{1}(\hat{\bf t},t_{l0})$ and thus both (43) and (45) fail. The case of NP-column requires an approach that leads to an expanded form of $C_{\bf s}^{1}$ contribution.
4.2.2 General form
To obtain a general form of the contribution $C_{\bf s}^{1}$ note that the relation $pc_{1l}+j_{l}=s_{l}$ restricts $j_{l}$ value to $j_{l}=s_{l}\bmod{c_{1l}}$ and the sum in (43) reduces to a single term
$C_{l}^{1}=A_{1,j_{l}}(\hat{\bf t})\hat{\bf t}^{(s_{l}-j_{l})\hat{\bf c}_{1}/c_{1l}},\quad j_{l}=s_{l}\bmod{c_{1l}},$ (48)
where $A_{1,j_{l}}(\hat{\bf t})$ is given by (38) and we have
$C_{\bf s}^{1}$ $=$ $C[\hat{\bf t}^{\hat{\bf s}}]\left(\Pi_{1}^{-1}(\hat{\bf t})C_{l}^{1}\right)=\!\!\sum_{\hat{\bf j}=\hat{\bf N}_{1}^{-}}^{\hat{\bf N}_{1}^{+}}\!\!a_{1,\hat{\bf j},j_{l}}C[\hat{\bf t}^{\hat{\bf s}}]\left(\Pi_{1}^{-1}(\hat{\bf t})\hat{\bf t}^{\hat{\bf j}+(s_{l}-j_{l})\hat{\bf c}_{1}/c_{1l}}\right)$ (49)
$=$ $\!\!\sum_{\hat{\bf j}=\hat{\bf N}_{1}^{-}}^{\hat{\bf N}_{1}^{+}}\!\!a_{1,\hat{\bf j},j_{l}}C[\hat{\bf t}^{\hat{\bf s}-\hat{\bf j}-(s_{l}-j_{l})\hat{\bf c}_{1}/c_{1l}}]\left(1/\Pi_{1}(\hat{\bf t})\right),\quad j_{l}=s_{l}\bmod{c_{1l}}.$
Consider a summand in (49) for some arbitrary value of column $\hat{\bf j}$, it reduces to
$C[\hat{\bf t}^{\hat{\bf s}-\hat{\bf j}-(s_{l}-j_{l})\hat{\bf c}_{1}/c_{1l}}]\left(1/\Pi_{1}(\hat{\bf t})\right)=W({\bf s}^{1}({\bf j}),{\bf D}_{1}),\quad{\bf s}^{1}({\bf j})=\hat{\bf s}-\hat{\bf j}-(s_{l}-j_{l})\hat{\bf c}_{1}/c_{1l},$
and write
$C_{\bf s}^{1}=\bar{W}^{l}_{1}=\!\!\sum_{\hat{\bf j}=\hat{\bf N}_{1}^{-}}^{\hat{\bf N}_{1}^{+}}\!\!a_{1,\hat{\bf j},j_{l}}W(\hat{\bf s}-\hat{\bf j}-(s_{l}-j_{l})\hat{\bf c}_{1}/c_{1l},{\bf D}_{1}),\quad j_{l}=s_{l}\bmod{c_{1l}}.$ (50)
It is convenient to rewrite $\bar{W}^{l}_{1}$ using the columns ${\bf N}_{1}^{\pm}$ employed in (40)
$\bar{W}^{l}_{1}=\!\!\sum_{{\bf j}={\bf N}_{1}^{-}}^{{\bf N}_{1}^{+}}\!\!a_{1,{\bf j}}W({\bf s}^{1}({\bf j}),{\bf D}_{1})\delta_{j_{l},s_{l}\bmod{c_{1l}}},\quad{\bf s}^{1}({\bf j})=\frac{{\bf s}^{1}-\hat{\bf j}c_{1l}+j_{l}\hat{\bf c}_{1}}{c_{1l}}.$ (51)
We note here that (51) in most cases is optional but it is compulsory for the NP-column with $g_{1}=\gcd({\bf c}_{1})>1$ and its special case of Z-column with $\hat{\bf c}_{1}=0,\ g_{1}=c_{1l}>1$.
For the NP-column we observe from (45) that all columns ${\bf c}^{1}_{j}$ of ${\bf D}_{1}$ are proportional to $g_{1}$. This means that the partition $W(\hat{\bf s}-\hat{\bf j}-(s_{l}-j_{l})\hat{\bf c}_{1}/c_{1l},{\bf D}_{1})$ in (50) should vanish when $g_{1}$ does not divide its first argument. It is obvious that with $j_{l}=s_{l}\bmod{c_{1l}}$ the last term $(s_{l}-j_{l})\hat{\bf c}_{1}/c_{1l}$ is proportional to $g_{1}$ and this means that the summands in (50) are nonzero only if a set of conditions $(\hat{\bf s}-\hat{\bf j})\bmod{g_{1}}=0$ is satisfied. In order to prevent application of general form (51) in further steps of the reduction algorithm it is recommended to use its equivalent version
$\bar{W}^{l}_{1}=\!\!\sum_{{\bf j}={\bf N}_{1}^{-}}^{{\bf N}_{1}^{+}}\!\!a_{1,{\bf j}}W({\bf s}^{1}({\bf j})/g_{1},{\bf D}_{1}/g_{1})\delta_{j_{l},s_{l}\bmod{c_{1l}}},\ (\hat{\bf s}-\hat{\bf j})\bmod{g_{1}}=0,\ {\bf s}^{1}({\bf j})=\hat{\bf s}-\hat{\bf j}-(s_{l}-j_{l})\hat{\bf c}_{1}/c_{1l}.$ (52)
Finally performing the removal of $K_{1}$ negative elements from the last row of the matrix ${\bf D}_{1}$ similar to one discussed in Section 4.2.1 we obtain another form of (52)
$\bar{W}^{l}_{1}=\!\!\sum_{{\bf j}={\bf N}_{1}^{-}}^{{\bf N}_{1}^{+}}\!\!(-1)^{K_{1}}a_{1,{\bf j}}H(\tilde{s}^{1}_{l}({\bf j}))W(\tilde{\bf s}^{1}({\bf j})/g_{1},\tilde{\bf D}_{1}/g_{1})\delta_{j_{l},s_{l}\bmod{c_{1l}}},$ (53)
$\tilde{\bf s}^{1}({\bf j})={\bf s}^{1}({\bf j})+\sum_{j=1}^{K_{1}}{\bf c}^{1}_{j},\quad\tilde{\bf D}_{1}=\{\tilde{\bf c}^{1}_{1},\ldots,\tilde{\bf c}^{1}_{m}\}),$
where $\tilde{\bf c}^{1}_{j}$ are defined in (47). The vector partition $W(\tilde{\bf s}^{1}({\bf j})/g_{1},\tilde{\bf D}_{1}/g_{1})$ is ready for the next reduction step.
For the Z-column we have $\hat{\bf c}_{1}={\bf 0}$ and the argument $\hat{\bf s}-\hat{\bf j}-(s_{l}-j_{l})\hat{\bf c}_{1}/c_{1l}$ in (50) reduces to $\hat{\bf s}-\hat{\bf j}$. All columns of ${\bf D}_{1}$ are proportional to $g_{1}=c_{1l}$ and we deduce that (51) turns into
$\bar{W}^{l}_{1}=\!\!\sum_{{\bf j}={\bf N}_{1}^{-}}^{{\bf N}_{1}^{+}}\!\!a_{1,{\bf j}}W((\hat{\bf s}-\hat{\bf j})/c_{1l},{\bf D}_{1}/c_{1l})\delta_{j_{l},s_{l}\bmod{c_{1l}}},\quad(\hat{\bf s}-\hat{\bf j})\bmod{c_{1l}}=0,$ (54)
that also should be converted to a form similar to (53).
4.3 Zero in last row of matrix ${\bf D}$
Consider evaluation of a $l$-tuple partition with $c_{1l}=0$. Rewrite the system ${\bf s}={\bf D\cdot x}$ as
${\bf s}={\bf c}_{1}x_{1}+\sum_{i=2}^{m}{\bf c}_{i}x_{i}\Rightarrow{\bf s}^{\prime}_{1}=\sum_{i=2}^{m}{\bf c}_{i}x_{i},\quad{\bf s}^{\prime}_{1}={\bf s}-{\bf c}_{1}x_{1}=\{s_{1}-c_{11}x_{1},s_{2}-c_{12}x_{1},\ldots,s_{l}\}^{T}.$ (55)
[…]
$\bm{r}^{\prime}_{1k}\cdot\bm{x}^{\prime}_{1}=s_{k}-c_{1k}x_{1},\ \ (1\leq k\leq l-1),\quad\bm{r}^{\prime}_{1l}\cdot\bm{x}^{\prime}_{1}=s_{l}.$
The matrix $\bm{D}^{\prime}_{1}=\{\bm{r}^{\prime}_{11},\bm{r}^{\prime}_{12},\ldots,\bm{r}^{\prime}_{1l}\}$ made of the rows $\bm{r}^{\prime}_{1k}$ together with the column ${\bf s}^{\prime}_{1}$ define a vector partition $W({\bf s}^{\prime}_{1},\bm{D}^{\prime}_{1})$ for each nonnegative value $x_{1}\geq 0$. This means that
$W({\bf s},{\bf D})=\sum_{x_{1}=0}^{\infty}W({\bf s}^{\prime}_{1},\bm{D}^{\prime}_{1}),$ (56)
[…]
Consider the elimination of the P-column ${\bf c}_{1}^{i},\ i>1$ from the matrix $\bm{D}^{\prime}_{1}$ in the partition $W({\bf s}^{\prime}_{1},\bm{D}^{\prime}_{1})$. It is represented by a single term
$W({\bf s}^{i}-x_{1}{\bf c}^{i}_{1},\bm{D}^{\prime}_{1i}),\quad\bm{D}^{\prime}_{1i}=\{{\bf c}^{i}_{2},{\bf c}^{i}_{3},\ldots,{\bf c}^{i}_{m}\},$ (57)
i.e., the matrix $\bm{D}^{\prime}_{1i}$ is obtained from the matrix ${\bf D}_{i}$ by dropping the first column ${\bf c}^{i}_{1}$.
The definition (6) of the generating function for the vector partition $W({\bf s},{\bf D})$ implies the recursion
$W({\bf s},{\bf D})-W({\bf s}-{\bf c}_{1},{\bf D})=W({\bf s},\{c_{2},c_{3},\ldots,c_{m}\}),$ (58)
leading to
$W({\bf s}^{i},{\bf D}_{i})-W({\bf s}^{i}-{\bf c}^{i}_{1},{\bf D}_{i})=W({\bf s}^{i},\bm{D}^{\prime}_{1i}).$ (59)
Use this relation to estimate the infinite sum of the terms (57) to obtain
$\sum_{x_{1}=0}^{\infty}W({\bf s}^{i}-x_{1}{\bf c}^{i}_{1},\bm{D}^{\prime}_{1i})=\sum_{x_{1}=0}^{\infty}[W({\bf s}^{i}-x_{1}{\bf c}^{i}_{1},{\bf D}_{i})-W({\bf s}^{i}-(x_{1}+1){\bf c}^{i}_{1},{\bf D}_{i})]=W({\bf s}^{i},{\bf D}_{i}),$ (60)
as there always exists a positive integer $k$ such that $W({\bf s}^{i}-k{\bf c}^{i}_{1},{\bf D}_{i})$ evaluates to zero. It is true if at least one element of ${\bf c}^{i}_{1}=c_{il}\hat{\bf c}_{1}$ is positive which is equivalent to the existence of at least one positive element in the column ${\bf c}_{1}$. If this condition fails, we apply the column sign inversion procedure (46) to the column ${\bf c}_{1}$ with the corresponding update of the column ${\bf s}\Rightarrow{\bf s}-{\bf c}_{1}$. Write the vector partition $W({\bf s},{\bf D})$ as
$W({\bf s},{\bf D})=C_{\bf s}^{1}+\sum_{i=2}^{m}W({\bf s}^{i},{\bf D}_{i}),$
where $C_{\bf s}^{1}$ denotes (yet unknown) contribution of the column ${\bf c}_{1}$. The relation (60) shows that each contribution $W({\bf s}^{i},{\bf D}_{i})$ for $2\leq i\leq m$ can be written as the infinite sum of $W({\bf s}^{i}-x_{1}{\bf c}^{i}_{1},\bm{D}^{\prime}_{1i})$. This leads to
$W({\bf s},{\bf D})=\sum_{x_{1}=0}^{\infty}W({\bf s}^{\prime}_{1},\bm{D}^{\prime}_{1})=\sum_{i=2}^{m}\sum_{x_{1}=0}^{\infty}W({\bf s}^{i}-x_{1}{\bf c}^{i}_{1},\bm{D}^{\prime}_{1i})=\sum_{i=2}^{m}W({\bf s}^{i},{\bf D}_{i}),$ (61)
and the contribution $C_{\bf s}^{1}$ vanishes. This result is used in the example discussed in Appendix B.
If the matrix $\bm{D}^{\prime}_{1}$ has another column ${\bf c}_{j}$ with last zero element we have two options – either the columns ${\bf c}_{1}$ and ${\bf c}_{j}$ are collinear or they are not. In the case of noncollinear columns the contributions $C_{\bf s}^{i}$ are independent of each other and thus we have $C_{\bf s}^{1}=C_{\bf s}^{j}=0$ while for the collinear columns one should use the method described in Section 5.
5 Vector partition with collinear columns
The reduction algorithm described in Section 4 cannot be applied in case when in the generator matrix ${\bf D}$ has collinear columns as it leads to appearance of the zero columns in the reduced partition.
Call an integer generator matrix competent if its columns (viewed as vectors) are noncollinear and incompetent otherwise. Assume that the vectors corresponding to the first $n$ columns of the incompetent generator matrix ${\bf D}$ are parallel to each other and rewrite the linear system ${\bf D}\cdot{\bf x}={\bf s}$ as follows
${\bf s}=\sum_{i=1}^{m}{\bf c}_{i}x_{i},\quad{\bf c}_{j}=u_{j}{\bf C},\quad{\bf C}=\{C_{1},C_{2},\ldots,C_{l}\}^{T},\quad 1\leq j\leq n<m,$ (62)
where $u_{j}$ are positive integers and vector ${\bf C}$ represents the GCD vector that divides elementwise all the collinear vectors ${\bf c}_{i},1\leq i\leq n<m$. The problem (62) is equivalent to
${\bf S}(k)={\bf s}-k{\bf C}=\sum_{i=n+1}^{m}{\bf c}_{i}x_{i},\quad k=\sum_{i=1}^{n}u_{i}x_{i},\quad 0\leq k\leq K=\min(\lfloor s_{r}/C_{r}\rfloor),\quad 1\leq r\leq l,$ (63)
where $K$ is computed using only positive $C_{r}>0$. Introduce a vector ${\bf u}=\{u_{1},u_{2},\ldots,u_{n}\}$ and a competent matrix with $(m-n)$ columns $\bar{\bf D}=\{{\bf c}_{n+1},\ldots,{\bf c}_{m}\},$ for which the corresponding partition $W({\bf S}(k),\bar{\bf D})$ admits the reduction either as the sum of $W_{i}^{l}$ or the mixture of $W^{l}_{i}$ and $\bar{W}^{l}_{i}$. The number of solutions of the system (63) for given value of $k$ is equal to a number of solutions of the first equation given by the partition $W({\bf S}(k),\bar{\bf D})$ multiplied by a number of solutions of the second equation, i.e., by a scalar partition $W(k,{\bf u})$. Then the vector partition $W({\bf s},{\bf D})$ is equivalent to a convolution
$W({\bf s},{\bf D})=\sum_{k=0}^{K}W(k,{\bf u})H({\bf s}-k{\bf C})W({\bf s}-k{\bf C},\bar{\bf D}),\quad K=\min(\lfloor s_{r}/C_{r}\rfloor),\quad 1\leq r\leq l,$ (64)
of the scalar partition and the vector partition $W({\bf s}-k{\bf C},\bar{\bf D})$ built on the competent matrix $\bar{\bf D}$ (it admits the reduction described in Section 4). Note that the function $H({\bf s}-k{\bf C})$ in (64) is applied only to the elements of the vector ${\bf s}-k{\bf C}$ corresponding to the rows of $\bar{\bf D}$ void of negative generators.
This approach can be extended to the case when the matrix ${\bf D}$ has $r>1$ groups of the collinear columns. Each $p$-th group is characterized by its GCD vector ${\bf C}_{p}$ and the $n_{p}$ integers $u_{p,i}$ making up the vector ${\bf u}_{p}=\{u_{p,1},u_{p,2},\ldots,u_{p,n_{p}}\}$. Introduce the vector ${\bf k}=\{k_{1},k_{2},\ldots,k_{r}\}$ and (63) turns into
${\bf S}({\bf k})={\bf s}-\sum_{p=1}^{r}k_{p}{\bf C}_{p}=\sum_{i=n+1}^{m}{\bf c}_{i}x_{i},\quad k_{p}=\sum_{i=1}^{n_{p}}u_{p,i}x_{i},\quad n=\sum_{p=1}^{r}n_{p}.$ (65)
The vector partition $W({\bf s},{\bf D})$ evaluates to the convolution
$W({\bf s},{\bf D})=\sum_{\bf k=0}^{\bf K}H({\bf S}({\bf k}))W({\bf S}({\bf k}),\bar{\bf D})\cdot\prod_{p=1}^{r}W(k_{p},{\bf u}_{p}),\ K_{p}=\min(\lfloor s_{i}/C_{p,i}\rfloor),\ \sum_{\bf k=0}^{\bf K}\equiv\sum_{k_{1}=0}^{K_{1}}\ldots\sum_{k_{r}=0}^{K_{r}}.$ (66)
Consider a special case $n=m-1$, i.e., when the competent matrix $\bar{\bf D}=\{{\bf c}\}$ has a single column ${\bf c}$. The vector partition $W({\bf S}({\bf k}),\{{\bf c}\})=1$ only when ${\bf S}({\bf k})=k_{0}{\bf c}$ for some nonnegative integer $k_{0}$, otherwise it vanishes. Thus we observe that $W({\bf S}({\bf k}),\bar{\bf D})=\delta_{{\bf S}({\bf k}),k_{0}{\bf c}}$ and we obtain using (66)
$W({\bf s},{\bf D})=\sum_{\bf k=0}^{\bf K}\sum_{k_{0}=0}^{K_{0}}\delta_{{\bf S}({\bf k}),k_{0}{\bf c}}H({\bf S}({\bf k}))\cdot\prod_{p=1}^{r}W(k_{p},{\bf u}_{p}),\quad K_{0}=\max(\lfloor s_{i}/c_{i}\rfloor),$ (67)
where again only positive $c_{i}$ are used for $K_{0}$ evaluation.
The extreme case arises for $n=m$, when each column of the matrix ${\bf D}$ belongs to one of $r$ groups of the collinear columns. Use ${\bf k}^{\prime}=\{k_{1},k_{2},\ldots,k_{r-1}\}$ to write
${\bf S}({\bf k}^{\prime})={\bf s}-\sum_{p=1}^{r-1}k_{p}{\bf C}_{p}=k_{r}{\bf C}_{r}=\sum_{i=1}^{n_{r}}u_{r,i}x_{i}{\bf C}_{r}.$ (68)
As ${\bf S}({\bf k}^{\prime})$ and ${\bf C}_{r}$ are collinear we obtain a set of additional conditions on $k_{r}$ leading to
$W({\bf s},{\bf D})=\sum_{{\bf k}=0}^{{\bf K}}\delta_{{\bf S}({\bf k}^{\prime}),k_{r}{\bf C}_{r}}H({\bf S}({\bf k}^{\prime}))\cdot\prod_{p=1}^{r}W(k_{p},{\bf u}_{p}),\quad K_{r}=\max(\lfloor s_{i}/C_{r,i}\rfloor).$ (69)
The example of partitions with collinear columns is discussed in Appendix C where we describe the reduction procedure for the matrix used in computation of the Kronecker coefficients [9, 10].
6 Iterative reduction of vector to scalar partitions
The basic step of iterative reduction procedure of the vector partition corresponding to the $l$-row matrix ${\bf E}$ described in Section 4 includes elimination of the last row and a single $i$-th column of ${\bf E}$. It results in a single partition based on the $(l-1)$-row matrix for the P-column or a superposition of partitions built on similar matrices for the NP-column. Each such matrix undergoes similar transformation in the consecutive iterative steps leading to a combination of scalar partitions. The total number of iterations required to reach this stage is $l-1$.
We note in Section 3.1 that the order of the rows in the two-row matrix ${\bf E}$ affects the individual scalar partition contributions to the double partition. The same holds for a matrix with the larger number of rows. This means that while one can change row order in the initial matrix ${\bf E}$, it is forbidden for any matrix generated in the course of the iterative elimination procedure. For definiteness assume that the order of row elimination is fixed and at every step we eliminate the last row of each generated matrix.
Consider a vector partition $A_{\nu}W({\bf E}_{\nu})$ corresponding to the $n\times(p+1)$ matrix ${\bf E}_{\nu}$ obtained after $\nu$ reduction steps of the original $l\times(m+1)$ matrix ${\bf E}$, so that $n=l-\nu$ and $p=n-\nu$. The factor $A_{\nu}$ is a product of the factors shown in (47) and (53) computed at each reduction step. The matrix ${\bf E}_{\nu}$ can be presented as a set of rows ${\bf E}_{\nu}=\{{\bf R}_{1},{\bf R}_{2},\ldots,{\bf R}_{n}\}$, where the row ${\bf R}_{k}=\{S_{k}=R_{k0},R_{k1},\ldots,R_{kp}\}$ has $p+1$ elements and $S_{k}$ is the linear superposition of the components $s_{i}$ of the original $l\times(m+1)$ matrix ${\bf E}$
$S_{k}=L^{\prime}_{k}({\bf s})=\kappa_{k0}+\sum_{j=1}^{m}\kappa_{kj}s_{j},$
with integer coefficients $\kappa_{kj}$. The last row ${\bf R}_{n}=\{S_{n}=R_{n0},R_{n1},\ldots,R_{np}\}$ has all nonnegative numerical elements while other rows ${\bf R}_{k}\ (1\leq k\leq n-1)$ might have negative elements.
How can one define the partition function in this case? The last row ${\bf R}_{n}$ of the matrix corresponds to a scalar partition defined by (1) that requires to find all sets ${\bf x}_{K}$ of nonnegative integer solutions. Assume that we find all such sets, their number is given by a scalar partition $W(S_{n},\{R_{n1},\ldots,R_{np}\})$. Now substitute each solution ${\bf x}_{K}$ into the remaining $n-1$ equations and retain only those that satisfy them all. The number of such solutions is the vector partition $W({\bf E}_{\nu})$. This definition fails when the last row has zero entries as the number of solutions ${\bf x}_{K}$ in this case is infinite. Nevertheless, one still can proceed with the evaluation procedure in these cases, this problem is addressed in Section 4.3 where we prove that the contribution of a column with the last zero element is zero.
In the original definition of the vector partition mentioned in Section 1 the elements of the vector ${\bf s}$ are considered nonnegative. This assumption does not hold in general for ${\bf E}_{\nu}$ as the linear superpositions $L^{\prime}_{k}({\bf s})$ might be negative for some values of the vector ${\bf s}$. For the rows with negative numerical elements the negative $L^{\prime}_{k}({\bf s})<0$ are acceptable. But if the row ${\bf R}_{k}$ is void of negative numbers we have to restrict its initial element $L^{\prime}_{k}({\bf s})$ to the nonnegative values $L^{\prime}_{k}({\bf s})\geq 0$ and to replace the partition $A_{\nu}W({\bf E}_{\nu})$ by
$A_{\nu}W({\bf E}_{\nu})\Rightarrow A_{\nu}B_{\nu}W({\bf E}_{\nu}),\quad B_{\nu}=\prod_{k}H(L^{\prime}_{k}({\bf s})).$ (70)
Application of the procedure (70) after every reduction step is required for determination of the partition chambers.
7 Conclusive remarks
In this manuscript we present an implementation of the Sylvester program aimed to reduction of vector (compound) partition to scalar partitions. The first step in this direction was made in 1860 by Cayley [4] when he showed that the double partition problem subject to a set of conditions admits a compact solution in which the $i$-th column ${\bf c}_{i}$ of a positive generator matrix ${\bf D}$ leads to a single scalar partition contribution $W_{i}^{2}$ (see Section 3.1). The author of this manuscript have suggested an alternative approach that solves the problem when the original Cayley method fails because the conditions are not satisfied [14] .
The results presented in the manuscript show that Sylvester strongly underestimated the number of scalar partitions required for computation of the vector partition – their minimal number in case of $l\times m$ positive matrix might significantly surpass $m!/(m-l+1)!$ which in its turn is $(l-1)!$ times larger than predicted by Sylvester in [17]. We address the case of collinear columns of the generator matrix in Section 5 and show that in this case a vector partition boils down to a convolution of scalar partitions.
Our approach is a generalization of the algorithms discussed in [14] and here we demonstrate that a vector partition always reduces iteratively either to a sum of scalar partitions themselves or their convolutions, and thus the Sylvester program is successfully implemented. Recalling that the scalar partitions can be computed as a superposition of the known polynomials [11] we arrive at the important statement — any vector partition can be expressed through the Bernoulli polynomials of higher order.
All known approaches to evaluation of vector partitions (see [1, 15, 3, 18]), their later modifications [2, 5, 8] and the software packages based of these methods successfully determine the chambers but fail to produce the explicit expressions for the vector partitions as they replace the original generating function by the signed sum of the simpler partial generating functions valid in specific chamber. This means that for each new argument column ${\bf s}$ one has to perform the Taylor expansion of the partial generating functions to extract the desired coefficient. Our method is the first one that permits the derivation of a single explicit formula for the vector partition which is valid for an arbitrary value of ${\bf s}$. The usage of this formula does not require the explicit knowledge of the chambers but nevertheless it implicitly contains their description.
Appendices
Appendix A Triple partition reduction example
Consider a reduction of a triple partition $W({\bf E})$ defined by a matrix ${\bf E}$
${\bf E}=\left(\begin{array}[]{ccccc}s_{1}&1&2&1&1\\ s_{2}&2&1&3&4\\ s_{3}&3&2&3&2\end{array}\right),$ (A1)
and perform the first reduction step by eliminating the last row of ${\bf E}$. We obtain
$W({\bf E})=\sum_{k=1}^{4}W({\bf E}_{k}),$ (A2)
with
${\bf E}_{1}=\left(\begin{array}[]{cccc}3s_{1}-s_{3}&4&0&1\\ 3s_{2}-2s_{3}&-1&3&8\end{array}\right),\quad{\bf E}_{2}=\left(\begin{array}[]{cccc}3s_{1}-s_{3}&0&4&1\\ 3s_{2}-3s_{3}&-3&-3&6\end{array}\right),$ (A7)
${\bf E}_{3}=\left(\begin{array}[]{cccc}2s_{1}-s_{3}&-1&2&-1\\ 2s_{2}-4s_{3}&-8&-6&-6\end{array}\right),\quad{\bf E}_{4}=\left(\begin{array}[]{cccc}2s_{1}-2s_{3}&-4&-4&-2\\ 2s_{2}-s_{3}&1&3&6\end{array}\right),$ (A12)
where each matrix ${\bf E}_{i}$ contains the NP-column. Perform the following transformation of the matrices ${\bf E}_{i}$ – first divide each row by the GCD of its numerical elements, and then use (46, 47) to make last row positive, and obtain a sum of four double partitions
$W({\bf E})=\sum_{k=1}^{4}a_{k}W(\tilde{\bf E}_{k}),\quad a_{1},a_{3}=-1,\quad a_{2},a_{4}=1,$ (A13)
with
$\tilde{\bf E}_{1}=\left(\begin{array}[]{cccc}3s_{1}-s_{3}+4&-4&0&1\\ 3s_{2}-2s_{3}-1&1&3&8\end{array}\right),$ $\tilde{\bf E}_{2}=\left(\begin{array}[]{cccc}3s_{1}-s_{3}+4&0&-4&1\\ s_{2}-s_{3}-2&1&1&2\end{array}\right),$ (A18)
$\tilde{\bf E}_{3}=\left(\begin{array}[]{cccc}2s_{1}-s_{3}&1&-2&1\\ s_{2}-2s_{3}-10&4&3&3\end{array}\right),$ $\tilde{\bf E}_{4}=\left(\begin{array}[]{cccc}s_{1}-s_{3}&-2&-2&-1\\ 2s_{2}-s_{3}&1&3&6\end{array}\right).$ (A23)
As discussed in Section 6 each summand in (A13) acquires an additional factor producing
$W({\bf E})=\sum_{k=1}^{4}A_{k}W(\tilde{\bf E}_{k}),$ (A24)
where
$A_{1}=-H(3s_{2}-2s_{3}-1),\quad A_{2}=H(s_{2}-s_{3}-2),\quad A_{3}=-H(s_{2}-2s_{3}-10),\quad A_{4}=H(2s_{2}-s_{3}).$ (A25)
Each double partition in (A24) to be reduced to a set of scalar partitions as shown in Section 3. Note that $\tilde{\bf E}_{1}$ has a single NP-column $\{0,3\}^{T}$ and thus, the elimination of this column should be done using the general Cayley algorithm presented in Section 3.2, while all other columns are processed as shown in Section 3.1. The results for $W(\tilde{\bf E}_{k})$ read
$W(\tilde{\bf E}_{1})$ $=$ $W(s_{1}+4s_{2}-3s_{3},\{4,11\})+W(8s_{1}-s_{2}-2s_{3}-1,\{1,11\})$
-- $\sum_{j=2}^{4}W((3s_{1}-s_{3}+4)/3-j,\{1,4\})\delta_{(3s_{2}-2s_{3}-1)\bmod{3},0}$
-- $\sum_{j=1,2,4}W((3s_{1}-s_{3}+2)/3-j,\{1,4\})\delta_{(3s_{2}-2s_{3}-2)\bmod{3},0}$
-- $\sum_{j=0,2,3}W((3s_{1}-s_{3})/3-j,\{1,4\})\delta_{(3s_{2}-2s_{3})\bmod{3},0}\;,$
$W(\tilde{\bf E}_{2})$ $=$ $W(3s_{1}+4s_{2}-5s_{3}-4,\{9,4\})-W(3s_{1}-s_{3},\{1,4\})+W(6s_{1}-s_{2}-s_{3}-4,\{1,9\}),$
$W(\tilde{\bf E}_{3})$ $=$ $W(6s_{1}+2s_{2}-7s_{3}-20,\{9,11\})-W(8s_{1}-s_{2}-2s_{3}-1,\{1,11\})$
$+$ $W(6s_{1}-s_{2}-s_{3},\{1,9\}),$
$W(\tilde{\bf E}_{4})$ $=$ $W(s_{1}+4s_{2}-3s_{3},\{4,11\})-W(3s_{1}+4s_{2}-5s_{3}-4,\{4,9\})$
$+$ $W(6s_{1}+2s_{2}-7s_{3}-20,\{9,11\}).$
These expressions together with (A24,A25) give the complete reduction of the triple partition defined by the matrix (A1) into a sum of twenty scalar partitions. Note that the first argument $L({\bf s})$ of each scalar partition should be nonnegative that together with the factor $A_{i}$ in (A24) completely determine the scalar partition chamber.
Appendix B Alternative triple partition reduction
In the Appendix A after the first reduction step we encounter a matrix $\tilde{\bf E}_{1}$ having NP-column $\{0,3\}^{T}$ that requires application of the general Cayley algorithm for its reduction. The same time in Section 4.3 we prove that the reduction of a column of the type $\{b,0\}^{T}$ with positive $b$ leads to vanishing contribution. It is instructive to reduce the number of terms in the final expression of the triple partition via the scalar ones. The convenient way to do that is to reshuffle the rows in the original triple matrix ${\bf E}$.
Consider reduction of a triple partition $W({\bf E}^{\prime})$ defined by the matrix
${\bf E}^{\prime}=\left(\begin{array}[]{ccccc}s_{2}&2&1&3&4\\ s_{1}&1&2&1&1\\ s_{3}&3&2&3&2\end{array}\right),$ (B1)
obtained from ${\bf E}$ by reordering of the first two rows. We find
$W({\bf E})=W({\bf E}^{\prime})=\sum_{k=1}^{4}a^{\prime}_{k}W(\tilde{\bf E}^{\prime}_{k}),\quad a^{\prime}_{1},a^{\prime}_{3}=-1,\quad a^{\prime}_{2},a^{\prime}_{4}=1,$ (B2)
where after the required transformations we obtain
$\tilde{\bf E}^{\prime}_{1}=\left(\begin{array}[]{cccc}2s_{2}-s_{3}&-1&-3&-6\\ s_{1}-s_{3}-5&2&2&1\end{array}\right),\quad\tilde{\bf E}^{\prime}_{2}=\left(\begin{array}[]{cccc}3s_{2}-2s_{3}&-1&8&3\\ 3s_{1}-s_{3}&4&1&0\end{array}\right),$ (B7)
$\tilde{\bf E}^{\prime}_{3}=\left(\begin{array}[]{cccc}s_{2}-s_{3}-1&1&-1&2\\ 3s_{1}-s_{3}&0&4&1\end{array}\right),\quad\tilde{\bf E}^{\prime}_{4}=\left(\begin{array}[]{cccc}s_{2}-2s_{3}-7&-3&4&3\\ 2s_{1}-s_{3}-2&2&1&1\end{array}\right).$ (B12)
Adding the $H$ factors we have
$W({\bf E}^{\prime})=\sum_{k=1}^{4}A^{\prime}_{k}W(\tilde{\bf E}^{\prime}_{k}),\ A^{\prime}_{1}=-H(s_{1}-s_{3}-5),\ A^{\prime}_{2}=-A^{\prime}_{3}=H(3s_{1}-s_{3}),\ A^{\prime}_{4}=H(2s_{1}-s_{3}-2).$
The second reduction step turns each double partition $W(\tilde{\bf E}^{\prime}_{k})$ into a sum of scalar partitions, but this time the contribution of the columns $\{3,0\}^{T}$ in $\tilde{\bf E}^{\prime}_{2}$ and $\{1,0\}^{T}$ in $\tilde{\bf E}^{\prime}_{3}$ vanishes. Thus we find
$W(\tilde{\bf E}^{\prime}_{1})$ $=$ $W(s_{1}+4s_{2}-3s_{3},\{4,11\})-W(3s_{1}+4s_{2}-5s_{3}-4,\{4,9\})$ (B13)
$+$ $W(6s_{1}+2s_{2}-7s_{3}-20,\{9,11\}),$
$W(\tilde{\bf E}^{\prime}_{2})$ $=$ $W(3s_{1}+4s_{2}-5s_{3},\{11,4\})-W(-8s_{1}+s_{2}+2s_{3}-11,\{1,11\}),$
$W(\tilde{\bf E}^{\prime}_{3})$ $=$ $W(3s_{1}+4s_{2}-5s_{3},\{9,4\})-W(-6s_{1}+s_{2}+s_{3}-10,\{1,9\}),$
$W(\tilde{\bf E}^{\prime}_{4})$ $=$ $W(6s_{1}+2s_{2}-7s_{3}-20,\{9,11\})+W(-8s_{1}+s_{2}+2s_{3}-11,\{1,11\})$
-- $W(-6s_{1}+s_{2}+s_{3}-10,\{1,9\}).$
The total number of scalar partitions in (B13) is ten which is significantly smaller than we have in Appendix A.
Appendix C Partition with collinear columns
Consider reduction of a triple partition $W({\bf E})$ defined by the matrix used in the computation of the Kronecker coefficients [9, 10]
${\bf E}=\left(\begin{array}[]{cccccccccccc}s_{3}&0&0&1&1&1&1&2&1&2&2&3\\ s_{2}&0&1&0&0&1&1&1&1&1&2&2\\ s_{1}&1&0&0&1&0&0&0&1&1&1&1\end{array}\right).$ (C1)
It has two collinear columns $\{1,1,0\}^{T}$ but their contribution vanishes due to last zero element. As the last row of ${\bf E}$ has five zeros the first reduction step produces only six nonvanishing terms
$W({\bf E})=\sum_{k=1}^{6}A_{6}W({\bf E}_{k}),\quad A_{1}=H(s_{2})H(s_{3}),$ $A_{2}=-H(s_{2}-2s_{1}-6)H(s_{3}-3s_{1}-9),$
$A_{3}=-H(s_{2})H(s_{3}-s_{1}-1),$ $A_{4}=H(s_{2}-2s_{1}-6)H(s_{3}-2s_{1}-4),$
$A_{5}=H(s_{2}-s_{1}-2)H(s_{3}-s_{1}-1),$ $A_{6}=-H(s_{2}-s_{1}-2)H(s_{3}-2s_{1}-4),$
and the matrices ${\bf E}_{k}$ read
${\bf E}_{1}=\left(\begin{array}[]{ccccc|cc|cc|cc}s_{3}&1&1&1&2&2&2&1&1&3&0\\ s_{2}&1&1&1&2&1&1&0&0&2&1\end{array}\right),$ (C5)
${\bf E}_{2}=\left(\begin{array}[]{ccccc|cc|cc|cc}s_{3}-3s_{1}-9&1&1&1&2&2&2&1&1&3&0\\ s_{2}-2s_{1}-6&1&1&1&2&1&1&0&0&2&1\end{array}\right),$ (C8)
${\bf E}_{3}=\left(\begin{array}[]{ccccc|cc|cc|cc}s_{3}-s_{1}-1&1&1&1&2&0&0&1&1&1&2\\ s_{2}&1&1&1&2&1&1&0&0&2&1\end{array}\right),$ (C11)
${\bf E}_{4}=\left(\begin{array}[]{ccccc|cc|cc|cc}s_{3}-2s_{1}-4&1&1&1&2&0&0&1&1&1&2\\ s_{2}-2s_{1}-6&1&1&1&2&1&1&0&0&2&1\end{array}\right),$ (C14)
${\bf E}_{5}=\left(\begin{array}[]{ccccc|cc|cc|cc|}s_{3}-s_{1}-1&1&1&1&1&2&2&1&1&0&0\\ s_{2}-s_{1}-2&1&1&1&1&1&1&0&0&1&1\end{array}\right),$ (C17)
${\bf E}_{6}=\left(\begin{array}[]{ccccc|cc|cc|cc|}s_{3}-2s_{1}-4&1&1&1&1&2&2&1&1&0&0\\ s_{2}-s_{1}-2&1&1&1&1&1&1&0&0&1&1\end{array}\right),$ (C20)
where the groups of collinear columns are bounded from the right by a vertical line. Recalling that $W({\bf E}_{k})\equiv W({\bf s}_{k},{\bf D}_{k})$ we find three pairs of matrices with coinciding numerical parts ${\bf D}_{k}$ and different argument columns ${\bf s}_{k}$ – namely, ${\bf D}_{k}={\bf D}_{k+1}$ for $k=1,3,5$.
Use the recipies from Section 5 to rewrite $W({\bf s}_{1},{\bf D}_{1})$. We have ${\bf C}_{1}=\{1,1\}^{T}$ with ${\bf u}_{1}=\{1,1,1,2\}$; ${\bf C}_{2}=\{2,1\}^{T}$ with ${\bf u}_{2}=\{1,1\}$, and ${\bf C}_{3}=\{1,0\}^{T}$ with ${\bf u}_{3}=\{1,1\}$. Use (65) to write
${\bf S}_{1}({\bf k})={\bf s}_{1}-\sum_{p=1}^{3}k_{p}{\bf C}_{p}=\{s_{3}-k_{1}-2k_{2}-k_{3},s_{2}-k_{1}-k_{2}\}^{T},\quad\bar{\bf D}_{1}=\left(\begin{array}[]{cc}3&0\\ 2&1\end{array}\right),$ (C21)
and
$\!\!\!W({\bf s}_{1},{\bf D}_{1})\!=\!\!\!\sum_{k_{1}=0}^{K_{1}}\sum_{k_{2}=0}^{K_{2}}\sum_{k_{3}=0}^{K_{3}}H({\bf S}_{1}({\bf k}))W({\bf S}_{1}({\bf k}),\bar{\bf D}_{1})W(k_{1},\{1,1,1,2\})W(k_{2},\{1,1\})W(k_{3},\{1,1\}).$ (C22)
The upper limits $K_{i}$ in (C22) are determined by the nonnegativeness of the elements of the vector ${\bf S}_{1}({\bf k})$ in (C21). The relation (C22) admits further simplification as the scalar partition of any integer number $k$ into a set of $m$ units equals the binomial $\binom{k+m-1}{m-1}$, providing $W(k_{i},\{1,1\})=\binom{k_{i}+1}{1}=k_{i}+1$. The reduction of the double partition $W({\bf S}_{1}({\bf k}),\bar{\bf D}_{1})$ leads to
$W({\bf S}_{1}({\bf k}),\bar{\bf D}_{1})=W(s_{3}-k_{1}-2k_{2}-k_{3},\{3\})-W(2s_{3}-3s_{2}+k_{1}-k_{2}-2k_{3}-3,\{3\}),$ (C23)
where both scalar partitions have the form $W(s,\{n\})$ that evaluates to $W(s,\{n\})=\delta_{s\bmod{n},0}$ for nonnegative $s$. Collecting these results we reduce (C22) to the following combination of the scalar partitions
$W({\bf s}_{1},{\bf D}_{1})$ $=$ $\sum_{k_{1}=0}^{K_{1}}\sum_{k_{2}=0}^{K_{2}}\sum_{k_{3}=0}^{K_{3}}H({\bf S}_{1}({\bf k}))(k_{2}+1)(k_{3}+1)W(k_{1},\{1,1,1,2\})$ (C24)
$\times$ $\left[W(s_{3}-k_{1}-2k_{2}-k_{3},\{3\})-W(2s_{3}-3s_{2}+k_{1}-k_{2}-2k_{3}-3,\{3\})\right].$
The term $W({\bf s}_{2},{\bf D}_{2})=W({\bf s}_{2},{\bf D}_{1})$ produces the relation similar to (C22) with the only change ${\bf S}_{1}({\bf k})\Rightarrow{\bf S}_{2}({\bf k})$ as follows
${\bf S}_{2}({\bf k})={\bf s}_{2}-\sum_{p=1}^{3}k_{p}{\bf C}_{p}=\{s_{3}-3s_{1}-9-k_{1}-2k_{2}-k_{3},s_{2}-2s_{1}-6-k_{1}-k_{2}\}^{T},$ (C25)
leading to
$W({\bf S}_{2}({\bf k}),\bar{\bf D}_{2})=W(s_{3}-3s_{1}-k_{1}-2k_{2}-k_{3},\{3\})-W(2s_{3}-3s_{2}+k_{1}-k_{2}-k_{3}-3,\{3\}).$ (C26)
Turn to evaluation of $W({\bf s}_{3},{\bf D}_{3})$ that is similar to $W({\bf s}_{1},{\bf D}_{1})$. We have ${\bf C}_{1}=\{1,1\}^{T}$ with ${\bf u}_{1}=\{1,1,1,2\}$; ${\bf C}_{2}=\{0,1\}^{T}$ with ${\bf u}_{2}=\{1,1\}$, and ${\bf C}_{3}=\{1,0\}^{T}$ with ${\bf u}_{3}=\{1,1\}$. Use (65) to write
${\bf S}_{3}({\bf k})={\bf s}_{3}-\sum_{p=1}^{3}k_{p}{\bf C}_{p}=\{s_{3}-s_{1}-1-k_{1}-k_{3},s_{2}-k_{1}-k_{2}\}^{T},$ (C27)
and
$W({\bf s}_{3},{\bf D}_{3})=\sum_{{\bf k}=0}^{\bf K}(k_{2}+1)(k_{3}+1)H({\bf S}_{3}({\bf k}))W({\bf S}_{3}({\bf k}),\bar{\bf D}_{3})W(k_{1},\{1,1,1,2\}),\ \bar{\bf D}_{3}=\left(\begin{array}[]{cc}1&2\\ 2&1\end{array}\right),$ (C28)
where the double partition $W({\bf S}_{3}({\bf k}),\bar{\bf D}_{3})$ reduces to
$W({\bf S}_{3}({\bf k}),\bar{\bf D}_{3})$ $=$ $W(2s_{3}-2s_{1}-s_{2}-2-k_{1}+k_{2}-2k_{3},\{3\})$ (C29)
-- $W(s_{3}-2s_{2}-s_{1}+k_{1}+2k_{2}-k_{3}-4,\{3\}).$
Similarly we obtain
$W({\bf S}_{4}({\bf k}),\bar{\bf D}_{4})$ $=$ $W(2s_{3}-2s_{1}-s_{2}-2-k_{1}+k_{2}-2k_{3},\{3\})$ (C30)
-- $W(s_{3}-2s_{2}+2s_{1}+k_{1}+2k_{2}-k_{3}+5,\{3\}),$
that replaces $W({\bf S}_{3}({\bf k}),\bar{\bf D}_{3})$ in (C28) to find explicit expression for $W({\bf s}_{4},{\bf D}_{4})$. Finally, consider $W({\bf s}_{5},{\bf D}_{5})$ with the matrix ${\bf D}_{5}$ void of noncollinear columns. This case is served by the relations (68,69) derived in Section 5. We choose the vectors ${\bf C}_{p}$ as follows ${\bf C}_{1}=\{2,1\}^{T}$ with ${\bf u}_{1}=\{1,1\}$; ${\bf C}_{2}=\{1,0\}^{T}$ with ${\bf u}_{2}=\{1,1\}$; ${\bf C}_{3}=\{0,1\}^{T}$ with ${\bf u}_{3}=\{1,1\}$, and ${\bf C}_{4}=\{1,1\}^{T}$ with ${\bf u}_{4}=\{1,1,1,1\}$. Use the vectors ${\bf s}_{5}=\{s_{3}-s_{1}-1,s_{2}-s_{1}-2\},\ {\bf s}_{6}=\{s_{3}-2s_{1}-4,s_{2}-s_{1}-2\}$ to find
${\bf S}_{5}({\bf k}^{\prime})={\bf s}_{5}-\sum_{p=1}^{3}k_{p}{\bf C}_{p}=\{s_{3}-s_{1}-1-2k_{1}-k_{2},s_{2}-s_{1}-2-k_{1}-k_{3}\}=\{k_{4},k_{4}\},$ (C31)
and the similar relation for ${\bf S}_{6}({\bf k}^{\prime})$. The expression (69) turns into
$W({\bf s}_{5},{\bf D}_{5})\!\!\!\!\!$ $=$ $\!\!\!\!\!\sum_{{\bf k}=0}^{{\bf K}}H({\bf S}_{5}({\bf k}^{\prime}))\binom{k_{4}+3}{3}\delta_{s_{3}-s_{1}-1-2k_{1}-k_{2},k_{4}}\delta_{s_{2}-s_{1}-2-k_{1}-k_{3},k_{4}}\prod_{j=1}^{3}(k_{j}+1),$ (C32)
$W({\bf s}_{6},{\bf D}_{6})\!\!\!\!\!$ $=$ $\!\!\!\!\!\sum_{{\bf k}=0}^{{\bf K}}H({\bf S}_{6}({\bf k}^{\prime}))\binom{k_{4}+3}{3}\delta_{s_{3}-2s_{1}-4-2k_{1}-k_{2},k_{4}}\delta_{s_{2}-s_{1}-2-k_{1}-k_{3},k_{4}}\prod_{j=1}^{3}(k_{j}+1).$ (C33)